# FREE-Edit: Using Editing-aware Injection in Rectified Flow Models for Zero-shot Image-Driven Video Editing

### 01 Mar 2026 v1

@@ 5.4 Quantitative Results
Besides, we record the time consumption of each method over 10 editing trials with 41 frames per edit. Thanks to the real-time performance of LTX-Video-2B [8], our FREE-Edit achieves the fastest speed. The three variants of our FREE-Edit achieve roughly comparable speeds. Note that I2VEdit requires additional training for motion extraction, where it takes about 2 hours for a 41-frame editing video on an NVIDIA A100 GPU.
@@
5.5 Comparison With Text-guided Methods
5.6 Extension to Different RF models
6 Limitations
@@ Appendix A More Visualization Results
Recall that in Fig. 4 in our main paper, we present a qualitative comparison between FREE-Edit w/o injection, w/ vanilla injection, and w/ REE injection on adding object and object replacement. In this Appendix, we provide additonal example on style transfer. Ap-Fig. 2 shows that FREE-Edit w/o injection struggles to preserve motion fidelity in the source video. Even in the style transfer case, vanilla injection tends to encounter unwanted semantic leakage (see the nose and ears of the rhino). In the last row, we show the editing masks warped by the first-frame mask.

### 22 Mar 2026 v2

@@ 5.4 Quantitative Results
Besides, we record the time consumption of each method over 10 editing trials with 41 frames per edit. Thanks to the real-time performance of LTX-Video-2B [9], our FREE-Edit achieves the fastest speed. The three variants of our FREE-Edit achieve roughly comparable speeds. Note that I2VEdit requires additional training for motion extraction and takes about 2 hours for a 41-frame editing video on an NVIDIA A100 GPU.
@@
5.5 Comparison With Text-guided Methods
5.6 Extension to Different RF models
6 Limitations
@@ Appendix A More Visualization Results
Recall that in Fig. 4 in our main paper, we present a qualitative comparison between FREE-Edit w/o injection, w/ vanilla injection, and w/ REE injection on adding object and object replacement. In this Appendix, we provide an additional example on style transfer. Ap-Fig 2 shows that FREE-Edit w/o injection struggles to preserve motion fidelity in the source video. Even in the style transfer case, vanilla injection tends to encounter unwanted semantic leakage (see the nose and ears of the rhino). In the last row, we show the editing masks warped by the first-frame mask.